The most recent financial statements for Shinoda Manufacturing Co. are shown below: Income Statement Balance Sheet Sales $ 64,700 Current assets $ 30,500 Debt $ 46,700 Costs 44,430 Fixed assets 83,400 Equity 67,200 Taxable income $ 20,270 Total $ 113,900 Total $ 113,900 Tax (40%) 8,108 Net Income $ 12,162 Assets and costs are proportional to sales. Debt and equity are not. The company maintains a constant 48 percent dividend payout ratio. No external financing is possible. What is the internal growth rate? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) Internal growth rate %
Answer #1

ROA=net income/Total assets

=(12162/113900)=0.106777875

Retention ratio=1-payout ratio

=(1-0.48)=0.52

Internal growth rate=(ROA*Retention ratio)/[1-(ROA*Retention ratio)]

=(0.106777875*0.52)[1-(0.106777875*0.52)]

=5.88%(Approx).